Hereafter are summarized some of the activities held during these last two weeks:
Gerado Avila and Bernard Delabre have been in Lisbon to make the optical PAE tests of the UT4 Coudé Train optics during week 33 and of the UT1 CT during week 34.
In Geneva, Matteo has remeasured the new cameras flanges and unfortunately the bad planarity is still of the same value. A new corrective action is being studied.
In Garching, the first tests of the blue camera have been done. The test setup appears good and globally gave good results, once a small problem modifying the focal distance, thus giving a very big spot size, was solved. Some fine-tuning is still to be done and further tests may take place.
At IAC, the Main collimator has been checked by Ana and Felix Gracia. It is acceptable as the vertex is in good position. Some marks have been detected on the surface, but may be due to test setup at the manufacturer, this will be noted in the acceptance report, and the gravity of the problem has to be evaluated before final acceptance. The surfaces where the coating was removed to prepare the gluing for the mounting have been inverted by the manufacturer. Felix corrected already this by properly grinding the correct places, manually with a adequate grinding tool of 75 um. The mirror and the mounting structure and tooling have been packed and sent to Geneva on August 23.